OpenStudy (anonymous):

A merry go round rotates 12 times per minute. Find: the angular velocity of the merry go round in radians per second.

OpenStudy (anonymous):

take the angle traced out in one minute, 12/60 of a complete circle, if 2pi is a full circle in rads, then 2pi/60 is the angle traced out in one minute, Angular velocity=change in angle/change in time

ganeshie8 (ganeshie8):

if \(\theta = 2 \pi \) then, u need to ask how much time time it takes for making that angle \(\theta\)